# Multilingual State Space Models for Structured Question Answering in Indic Languages

### v1

@@ Abstract
The diversity and complexity of Indic languages present unique challenges for natural language processing (NLP) tasks, particularly in the domain of question answering (QA).To address these challenges, this paper explores the application of State Space Models (SSMs) to build efficient and contextually aware QA systems tailored for Indic languages. SSMs are particularly suited for this task due to their ability to model long-term and short-term dependencies in sequential data, making them well-equipped to handle the rich morphology, complex syntax, and contextual intricacies characteristic of Indian languages. We evaluated multiple SSM architectures across diverse datasets representing various Indic languages and conducted a comparative analysis of their performance. Our results demonstrate that these models effectively capture linguistic subtleties, leading to significant improvements in question interpretation, context alignment, and answer generation. This work represents the first application of SSMs to question answering tasks in Indic languages, establishing a foundational benchmark for future research in this domain. Furthermore, we propose enhancements to existing SSM frameworks, optimizing their applicability to low-resource settings and multilingual scenarios prevalent in Indic languages.
Machine Learning, ICML
@@ 1 Introduction
In recent years, the field of Natural Language Processing (NLP) has witnessed significant advancements with the development of models capable of handling complex linguistic structures and long-range dependencies. State Space Models (SSMs) (Goel et al., 2022) have emerged as a promising alternative to traditional architectures like Transformers  (Vaswani et al., 2017), particularly in tasks requiring efficient long-sequence modeling. Models such as Mamba  (Gu & Dao, 2024) and its variants have demonstrated the ability to manage long-context language modeling with constant memory usage, addressing some limitations inherent in Transformer-based models  (Nguyen et al., 2024)  (Bourdois, 2022). Despite these advancements, the development of question-answering (QA) systems for low-resource languages, particularly Indic languages, remains a challenging endeavor. India has a population of 1.4 billion and is home to 122 languages and 270 mother tongues. Indian languages fall in the minority as far as NLP models are concerned. One of the Indian languages—Hindi—is spoken by 577.8 million people worldwide  (Sabane et al., 2024). Likewise, Marathi, another Indian language, is ranked 11th on the list of most spoken languages. However, the scarcity of annotated datasets and neural models has hindered progress in building effective QA systems for these languages. Recent efforts have been made to address this gap by introducing large-scale QA datasets for Indic languages, such as the Indic QA dataset  (Singh et al., 2024), which provides a substantial resource for developing systems tailored to the linguistic nuances of Indic languages. This paper explores the various SSM models, to build efficient and contextually aware QA systems for Indic languages. SSMs are particularly suited for this task due to their ability to model both long-term and short-term dependencies in sequential data, making them well-equipped to handle the rich morphology, complex syntax, and contextual intricacies characteristic of Indian languages. The main contributions of this paper are:
@@ 2.1 Advances in State Space Models
Traditional SSMs, such as the Kalman Filter  (Kalman, 1960) and Hidden Markov Models (HMMs)  (Rabiner & Juang, 1986), have been widely used for decades. While effective for many tasks, these models struggle with long-range dependencies and high-dimensional data  (Bourdois, 2022). To overcome these limitations, SSMs have emerged, combining the foundational principles of traditional SSMs with the expressive power of deep learning  (Sarrof et al., 2024). These approaches allow for scalability and adaptability in modern machine learning applications  (Goel et al., 2022).SSMs have emerged as a robust alternative to transformers in sequence modeling, excelling particularly in areas where transformers face inherent limitations  (Patro et al., 2024). Mathematically, an SSM can be represented as follows:
@@
The  Equation 1  (Gu et al., 2022) describes a framework where the state vector $\mathbf{x}_{t}$ at time $t$ represents latent variables encapsulating the system’s underlying dynamics. The state evolution from $t$ to $t+1$ is governed by the state transition matrix $\mathbf{A}$, while the input matrix $\mathbf{B}$ determines the effect of the control input $\mathbf{u}_{t}$ on the state transitions. The control input $\mathbf{u}_{t}$ serves as an external influence, allowing modifications to the state dynamics at time $t$. Additionally, process noise $\mathbf{w}_{t}$ is modeled as a Gaussian random variable with zero mean ($\mathcal{N}(0,\mathbf{Q})$) and covariance matrix $\mathbf{Q}$, capturing uncertainties and randomness in the state transitions. The covariance matrix $\mathbf{Q}$ quantifies the variability of the process noise $\mathbf{w}_{t}$, ensuring the model accounts for inherent stochasticity.
@@
Transformers rely on self-attention mechanisms to capture relationships across sequences  (Vaswani et al., 2017). While effective, their quadratic complexity with respect to sequence length leads to significant computational overhead and memory usage for long contexts  (Taha, 2025). In contrast, SSMs are designed with a focus on efficiency and scalability, leveraging state transitions and compact latent representations to model temporal dependencies  (Shakhadri et al., 2025). This design enables SSMs to process sequences with linear computational complexity, making them more efficient for handling long-range dependencies in applications such as natural language processing, time-series forecasting, and audio processing  (Liu et al., 2024). SSMs have emerged as a significant development in sequence modeling, providing an effective approach to capture temporal dynamics in various tasks  (Abreu et al., 2024). SSMs are grounded in their ability to model dynamic systems through latent state representations, which evolve over time according to predefined transition dynamics and are observed through noisy measurements  (Wang et al., 2024).Recent advancements in SSMs have led to the creation of hybrid architectures and enhanced methodologies that address key limitations in sequence modeling. For instance,Mamba  (Gu & Dao, 2023) (Linear-Time Sequence Modeling with Selective State Spaces) (Gu & Dao, 2024) introduces a mechanism where the parameters of the SSM are functions of the input, allowing selective propagation or forgetting of information. This selective capability improves the model’s handling of discrete data and enables linear scaling with sequence length, outperforming traditional transformer models of similar size across diverse modalities such as language and audio. Another contribution is the Samba  (Ren et al., 2024), a hybrid architecture that fuses selective SSMs with sliding window attention, enabling efficient handling of unlimited context. By compressing sequences into recurrent hidden states while leveraging attention for precise recall, Samba demonstrates a remarkable ability to manage long sequences with minimal computational overhead. Similarly, models like Hymba  (Dong et al., 2024) and Jamba  (Lieber et al., 2024) integrate SSMs with attention mechanisms within hybrid architectures to achieve a balance between efficient summarization of long-range contexts and high-resolution recall. These models exemplify how combining the strengths of SSMs and transformers  (Vaswani et al., 2017) can yield state-of-the-art results, especially in tasks requiring both scalability and accuracy. Another critical area where SSMs surpass transformers is in inference speed and memory efficiency. Attention-free SSM models like Falcon Mamba  (Zuo et al., 2024) have proven to be significantly faster during inference while being lighter in memory usage for processing long sequences. For example, Falcon Mamba can match or exceed the performance of leading transformer-based models, such as Mistral 7B  (Jiang et al., 2023b) and Llama  (Touvron et al., 2023b), while requiring far fewer computational resources. This makes SSMs particularly suitable for deployment in resource-constrained environments or applications where real-time processing is critical. Additionally, transformers often struggle with handling highly structured or continuous data such as audio and sensor streams. In these domains, the state-transition mechanisms of SSMs enable them to natively and effectively capture temporal dynamics, offering a clear advantage over transformers, which often require additional architectural modifications to achieve comparable results.
@@ 2.2 Indic Language Question Answering
Indic languages, with their diverse scripts, complex grammar, and linguistic variations, present unique challenges for QA systems  (Dani & Sathe, 2024). Features such as free word order, inflectional patterns, and compound words complicate NLP tasks. While initiatives like the IndicQA dataset  (Sabane et al., 2024) have advanced question answering in these languages, the scarcity of annotated datasets and neural models remains a significant challenge. Techniques such as transfer learning and domain adaptation have shown promise, with multilingual models fine-tuned on task-specific data achieving notable improvements in Hindi and Marathi QA tasks  (Jin et al., 2022).
Adopting the SQuAD format  (Rajpurkar et al., 2016), a widely used benchmark in QA research, facilitates the creation of structured datasets for Indic languages. This format includes context passages, questions, and corresponding answers with exact character positions, ensuring consistency and compatibility across QA frameworks. Translating and adapting the SQuAD format for Indic languages has proven effective, as demonstrated by the performance improvements of the IndicQA dataset when aligned with this structure  (Singh et al., 2024).
To address data scarcity, data augmentation techniques such as back-translation have been used to generate synthetic SQuAD-style datasets  (Khan et al., 2024). Fine-tuning multilingual pre-trained models like XLM-R  (Conneau et al., 2020) and mT5  (Xue et al., 2021) on such datasets has further improved performance in Indic QA tasks . The structured nature of the SQuAD format allows for robust evaluation of fact-based, reasoning-based, and opinion-based questions, making it particularly suitable for Indic languages  (Upadhyay et al., 2024). Aligning Indic-language datasets with this format enhances model performance and facilitates better comparisons across multilingual QA research.
@@ 3 Methodology
Our methodology focuses on developing a question-answering framework tailored for Indic languages by leveraging the capabilities of SSMs. Specifically, we experimented with various models, including Mamba, Mamba-2,Falcon Mamba, Jamba, Zamba, Samba, and Hymba, to address the linguistic diversity and complexity of Indic languages. Each model was evaluated for its ability to process questions, answers, and context efficiently while preserving the grammatical and semantic nuances of the respective languages. The overall framework for this system is illustrated in  Figure 1.
@@ 3.1.1 Tokenization
For our data preprocessing pipeline, we employed IndicNLPtokenizer  (Kunchukuttan, 2020), which is specifically designed to address the unique grammatical structures and linguistic characteristics of Indic languages. This tokenizer is particularly effective for handling the complex and diverse scripts of languages such as Hindi and Marathi. By maintaining consistent tokenization across the context (context about the question), question, and answer in our dataset, it enabled uniformity and alignment throughout the preprocessing pipeline. The tokenizer was trained on our corpus to capture language-specific patterns and generate a custom vocabulary tailored to Indic languages. This vocabulary included frequently occurring tokens and special tokens such as padding (<pad>), unknown tokens (<unk>), start-of-sentence (<sos>), and end-of-sentence (eos) markers, which are essential for enabling sequence alignment, handling missing or unknown words, and defining sentence boundaries during both training and inference. These functionalities ensure that the model can process variable-length inputs, manage incomplete data, and correctly interpret input sequences. After constructing the vocabulary, we performed encoding, where the starting position (answer_start) of the answer, representing the character index of the answer within the context, was mapped to its corresponding tokenized representation. This process generated the start (token_start) and end (token_end) token positions, which define the answer span within the tokenized sequence. These fields are essential for enabling the model to accurately learn and predict answer spans, ensuring precise alignment between the original text and its tokenized format for effective training and evaluation. The overall pipeline is illustrated in  Figure 2.
@@ 3.2 Training
Subsequently, we fine-tuned the SSM model for question answering in Indic languages, we designed a training pipeline to enhance the performance for question answering in Indic languages. The fine-tuning process employed a resource-efficient strategy using low-rank adaptation (LoRA) (Hu et al., 2021). LoRA employs a strategy of updating only the parameters in the projection layers of attention modules and the embeddings layer, leaving the rest of the model unchanged. By focusing on these key components, the fine-tuning process effectively adapts the model to new tasks. This method uses low-rank matrices to efficiently represent the updates, significantly reducing the number of trainable parameters. Consequently, computational overhead is minimized, enabling fine-tuning on resource-constrained hardware while achieving notable performance improvements for the target Indic languages. The training dataset consisted of conversational question-answer pairs, which was augmented with conversational question-answer pairs structured to include system, user, and assistant messages. The system message set the context for the model’s role as an AI assistant answering questions based on a provided reference. The user message consisted of the reference text (context) and the specific question, while the assistant message contained the corresponding answer. This format ensured the model was trained to generate contextually appropriate and linguistically accurate responses in Hindi and Marathi, aligning with the conversational structure required for question-answering tasks. The implementation details of the fine-tuning process, including hyperparameters and training configurations, are discussed later in  Section 4.2
@@ 3.3 Models
Falcon Mamba: Falcon Mamba 7B, is a pure Mamba-based large language model trained on 5.8 trillion tokens.Falcon Mamba 7B surpasses leading open-weight Transformer-based models, including Mistral 7B  (Jiang et al., 2023a), Llama  (Touvron et al., 2023a).Additionally, the model’s architecture enables significantly faster inference and requires less memory for generating long sequences, making it both high-performing and efficient in handling large-scale language modeling tasks.
Jamba: Jamba is a hybrid language model that interleaves Transformer and Mamba layers, combining the strengths of both architectures. Incorporating a mixture-of-experts (MoE)  (Lo et al., 2024) approach, Jamba increases model capacity while maintaining manageable active parameter usage. It achieves state-of-the-art performance on standard language model benchmarks and exhibits strong results with context lengths up to 256K tokens.
@@
Samba: A simple hybrid architecture that layer-wise combines Mamba with Sliding Window Attention (SWA)  (Beltagy et al., 2020). SAMBA  (Ren et al., 2024) selectively compresses sequences into recurrent hidden states while maintaining the ability to recall recent memories precisely. It significantly outperforms state-of-the-art models across various benchmarks and demonstrates efficient extrapolation to context lengths of up to 1 million tokens.

### v2

@@ Abstract
The diversity and complexity of Indic languages present unique challenges for natural language processing (NLP) tasks, particularly in the domain of question answering (QA).To address these challenges, this paper explores the application of State Space Models (SSMs) to build efficient and contextually aware QA systems tailored for Indic languages. SSMs are particularly suited for this task due to their ability to model long-term and short-term dependencies in sequential data, making them well-equipped to handle the rich morphology, complex syntax, and contextual intricacies characteristic of Indian languages. We evaluated multiple SSM architectures across diverse datasets representing various Indic languages and conducted a comparative analysis of their performance. Our results demonstrate that these models effectively capture linguistic subtleties, leading to significant improvements in question interpretation, context alignment, and answer generation. This work represents the first application of SSMs to question-answering tasks in Indic languages, establishing a foundational benchmark for future research in this domain. Furthermore, we propose enhancements to existing SSM frameworks, optimizing their applicability to low-resource settings and multilingual scenarios prevalent in Indic languages.
Multilingual State Space Models for Structured Question Answering in Indic Languages
Arpita Vats1thanks: Equal contribution., Rahul Raja211footnotemark: 1, Mrinal Mathur311footnotemark: 1, Vinija Jain4, Aman Chadha4 1Santa Clara University, 2Carnegie Mellon University, 3Georgia Institute of Technology, 4Stanford University
@@ 1 Introduction
In recent years, the field of Natural Language Processing (NLP) has witnessed significant advancements with the development of models capable of handling complex linguistic structures and long-range dependencies. State Space Models (SSMs) Goel et al. (2022) have emerged as a promising alternative to traditional architectures like Transformers  Vaswani et al. (2017), particularly in tasks requiring efficient long-sequence modeling. Models such as Mamba  Gu and Dao (2024) and its variants have demonstrated the ability to manage long-context language modeling with constant memory usage, addressing some limitations inherent in Transformer-based models  Nguyen et al. (2024)  Bourdois (2022). Despite these advancements, the development of question-answering (QA) systems for low-resource languages, particularly Indic languages, remains a challenging endeavor. India has a population of 1.4 billion and is home to 122 languages and 270 mother tongues. Indian languages fall in the minority as far as NLP models are concerned. One of the Indian languages—Hindi—is spoken by 577.8 million people worldwide  Sabane et al. (2024). Likewise, Marathi, another Indian language, is ranked 11th on the list of most spoken languages. However, the scarcity of annotated datasets and neural models has hindered progress in building effective QA systems for these languages. Recent efforts have been made to address this gap by introducing large-scale QA datasets for Indic languages, such as the Indic QA dataset  Singh et al. (2024), which provides a substantial resource for developing systems tailored to the linguistic nuances of Indic languages. This paper explores the various SSM models to build efficient and contextually aware QA systems for Indic languages. SSMs are particularly suited for this task due to their ability to model both long-term and short-term dependencies in sequential data, making them well-equipped to handle the rich morphology, complex syntax, and contextual intricacies characteristic of Indian languages. The main contributions of this paper are:
@@ 2.1 Advances in State Space Models
Traditional SSMs, such as the Kalman Filter  Kalman (1960) and Hidden Markov Models (HMMs)  Rabiner and Juang (1986), have been widely used for decades. While effective for many tasks, these models struggle with long-range dependencies and high-dimensional data  Bourdois (2022). To overcome these limitations, SSMs have emerged, combining the foundational principles of traditional SSMs with the expressive power of deep learning  Sarrof et al. (2024). These approaches allow for scalability and adaptability in modern machine learning applications  Goel et al. (2022).SSMs have emerged as a robust alternative to transformers in sequence modeling, excelling particularly in areas where transformers face inherent limitations  Patro et al. (2024). Mathematically, an SSM can be represented as follows:
@@
The  eq. 1  Gu et al. (2022) describes a framework where the state vector $\mathbf{x}_{t}$ at time $t$ represents latent variables encapsulating the system’s underlying dynamics. The state evolution from $t$ to $t+1$ is governed by the state transition matrix $\mathbf{A}$, while the input matrix $\mathbf{B}$ determines the effect of the control input $\mathbf{u}_{t}$ on the state transitions. The control input $\mathbf{u}_{t}$ serves as an external influence, allowing modifications to the state dynamics at time $t$. Additionally, process noise $\mathbf{w}_{t}$ is modeled as a Gaussian random variable with zero mean ($\mathcal{N}(0,\mathbf{Q})$) and covariance matrix $\mathbf{Q}$, capturing uncertainties and randomness in the state transitions. The covariance matrix $\mathbf{Q}$ quantifies the variability of the process noise $\mathbf{w}_{t}$, ensuring the model accounts for inherent stochasticity.
@@
Transformers rely on self-attention mechanisms to capture relationships across sequences  Vaswani et al. (2017). While effective, their quadratic complexity with respect to sequence length leads to significant computational overhead and memory usage for long contexts  Taha (2025). In contrast, SSMs are designed with a focus on efficiency and scalability, leveraging state transitions and compact latent representations to model temporal dependencies  Shakhadri et al. (2025). This design enables SSMs to process sequences with linear computational complexity, making them more efficient for handling long-range dependencies in applications such as natural language processing, time-series forecasting, and audio processing  Liu et al. (2024). SSMs have emerged as a significant development in sequence modeling, providing an effective approach to capture temporal dynamics in various tasks  Abreu et al. (2024). SSMs are grounded in their ability to model dynamic systems through latent state representations, which evolve over time according to predefined transition dynamics and are observed through noisy measurements  Wang et al. (2024).Recent advancements in SSMs have led to the creation of hybrid architectures and enhanced methodologies that address key limitations in sequence modeling. For instance, Mamba  Gu and Dao (2023) (Linear-Time Sequence Modeling with Selective State Spaces) Gu and Dao (2024) introduces a mechanism where the parameters of the SSM are functions of the input, allowing selective propagation or forgetting of information. This selective capability improves the model’s handling of discrete data and enables linear scaling with sequence length, outperforming traditional transformer models of similar size across diverse modalities such as language and audio. Another contribution is the Samba  Ren et al. (2024), a hybrid architecture that fuses selective SSMs with sliding window attention, enabling efficient handling of unlimited context. By compressing sequences into recurrent hidden states while leveraging attention for precise recall, Samba demonstrates a remarkable ability to manage long sequences with minimal computational overhead. Similarly, models like Hymba  Dong et al. (2024) and Jamba  Lieber et al. (2024) integrate SSMs with attention mechanisms within hybrid architectures to achieve a balance between efficient summarization of long-range contexts and high-resolution recall. These models exemplify how combining the strengths of SSMs and transformers  Vaswani et al. (2017) can yield state-of-the-art results, especially in tasks requiring both scalability and accuracy. Another critical area where SSMs surpass transformers is in inference speed and memory efficiency. Attention-free SSM models like Falcon Mamba  Zuo et al. (2024) have proven to be significantly faster during inference while being lighter in memory usage for processing long sequences. For example, Falcon Mamba can match or exceed the performance of leading transformer-based models, such as Mistral 7B  Jiang et al. (2023b) and Llama  Touvron et al. (2023b), while requiring far fewer computational resources. This makes SSMs particularly suitable for deployment in resource-constrained environments or applications where real-time processing is critical. Additionally, transformers often struggle with handling highly structured or continuous data such as audio and sensor streams. In these domains, the state-transition mechanisms of SSMs enable them to natively and effectively capture temporal dynamics, offering a clear advantage over transformers, which often require additional architectural modifications to achieve comparable results.
@@ 2.2 Indic Language Question Answering
Indic languages, with their diverse scripts, complex grammar, and linguistic variations, present unique challenges for QA systems  Dani and Sathe (2024). Features such as free word order, inflectional patterns, and compound words complicate NLP tasks. While initiatives like the IndicQA dataset  Sabane et al. (2024) have advanced question answering in these languages, the scarcity of annotated datasets and neural models remains a significant challenge. Techniques such as transfer learning and domain adaptation have shown promise, with multilingual models fine-tuned on task-specific data achieving notable improvements in Hindi and Marathi QA tasks  Jin et al. (2022).
Adopting the SQuAD format  Rajpurkar et al. (2016), a widely used benchmark in QA research, facilitates the creation of structured datasets for Indic languages. This format includes context passages, questions, and corresponding answers with exact character positions, ensuring consistency and compatibility across QA frameworks. Translating and adapting the SQuAD format for Indic languages has proven effective, as demonstrated by the performance improvements of the IndicQA dataset when aligned with this structure  Singh et al. (2024).
To address data scarcity, data augmentation techniques such as back-translation have been used to generate synthetic SQuAD-style datasets  Khan et al. (2024). Fine-tuning multilingual pre-trained models like XLM-R  Conneau et al. (2020) and mT5  Xue et al. (2021) on such datasets has further improved performance in Indic QA tasks . The structured nature of the SQuAD format allows for robust evaluation of fact-based, reasoning-based, and opinion-based questions, making it particularly suitable for Indic languages  Upadhyay et al. (2024). Aligning Indic-language datasets with this format enhances model performance and facilitates better comparisons across multilingual QA research.
@@ 3 Methodology
Our methodology focuses on developing a question-answering framework tailored for Indic languages by leveraging the capabilities of SSMs. Specifically, we experimented with various models, including Mamba, Mamba-2,Falcon Mamba, Jamba, Zamba, Samba, and Hymba, to address the linguistic diversity and complexity of Indic languages. Each model was evaluated for its ability to process questions, answers, and context efficiently while preserving the grammatical and semantic nuances of the respective languages. The overall framework for this system is illustrated in  fig. 1.
@@ 3.1.1 Tokenization
For our data preprocessing pipeline, we employed IndicNLPtokenizer  Kunchukuttan (2020), which is specifically designed to address the unique grammatical structures and linguistic characteristics of Indic languages. This tokenizer is particularly effective for handling the complex and diverse scripts of languages such as Hindi and Marathi. By maintaining consistent tokenization across the context (context about the question), question, and answer in our dataset, it enabled uniformity and alignment throughout the preprocessing pipeline. The tokenizer was trained on our corpus to capture language-specific patterns and generate a custom vocabulary tailored to Indic languages. This vocabulary included frequently occurring tokens and special tokens such as padding (<pad>), unknown tokens (<unk>), start-of-sentence (<sos>), and end-of-sentence (eos) markers, which are essential for enabling sequence alignment, handling missing or unknown words, and defining sentence boundaries during both training and inference. These functionalities ensure that the model can process variable-length inputs, manage incomplete data, and correctly interpret input sequences. After constructing the vocabulary, we performed encoding, where the starting position (answer_start) of the answer, representing the character index of the answer within the context, was mapped to its corresponding tokenized representation. This process generated the start (token_start) and end (token_end) token positions, which define the answer span within the tokenized sequence. These fields are essential for enabling the model to accurately learn and predict answer spans, ensuring precise alignment between the original text and its tokenized format for effective training and evaluation. The overall pipeline is illustrated in  fig. 2.
@@ 3.2 Training
Subsequently, we fine-tuned the SSM model for question answering in Indic languages, we designed a training pipeline to enhance the performance for question answering in Indic languages. The fine-tuning process employed a resource-efficient strategy using low-rank adaptation (LoRA) Hu et al. (2021). LoRA employs a strategy of updating only the parameters in the projection layers of attention modules and the embeddings layer, leaving the rest of the model unchanged. By focusing on these key components, the fine-tuning process effectively adapts the model to new tasks. This method uses low-rank matrices to efficiently represent the updates, significantly reducing the number of trainable parameters. Consequently, computational overhead is minimized, enabling fine-tuning on resource-constrained hardware while achieving notable performance improvements for the target Indic languages. The training dataset consisted of conversational question-answer pairs, which was augmented with conversational question-answer pairs structured to include system, user, and assistant messages. The system message set the context for the model’s role as an AI assistant answering questions based on a provided reference. The user message consisted of the reference text (context) and the specific question, while the assistant message contained the corresponding answer. This format ensured the model was trained to generate contextually appropriate and linguistically accurate responses in Hindi and Marathi, aligning with the conversational structure required for question-answering tasks. The implementation details of the fine-tuning process, including hyperparameters and training configurations, are discussed later in  section 4.2
@@ 3.3 Models
Falcon Mamba: Falcon Mamba 7B, is a pure Mamba-based large language model trained on 5.8 trillion tokens. Falcon Mamba 7B surpasses leading open-weight Transformer-based models, including Mistral 7B  Jiang et al. (2023a), Llama  Touvron et al. (2023a).Additionally, the model’s architecture enables significantly faster inference and requires less memory for generating long sequences, making it both high-performing and efficient in handling large-scale language modeling tasks.
Jamba: Jamba is a hybrid language model that interleaves Transformer and Mamba layers, combining the strengths of both architectures. Incorporating a mixture-of-experts (MoE)  Shazeer et al. (2017) approach, Jamba increases model capacity while maintaining manageable active parameter usage. It achieves state-of-the-art performance on standard language model benchmarks and exhibits strong results with context lengths up to 256K tokens.
@@
Samba: A simple hybrid architecture that layer-wise combines Mamba with Sliding Window Attention (SWA)  Beltagy et al. (2020). SAMBA  Ren et al. (2024) selectively compresses sequences into recurrent hidden states while maintaining the ability to recall recent memories precisely. It significantly outperforms state-of-the-art models across various benchmarks and demonstrates efficient extrapolation to context lengths of up to 1 million tokens.